# Rigid clusters in shear-thickening suspensions: a nonequilibrium critical transition

### v3

@@ Acknowledgments
This work was supported by NSF CBET-2228681, CBET-2228680, and DMR-2026834. AS acknowledges Aryabhata HPC facility of IIT (ISM) Dhanbad and National Supercomputing Mission (NSM) for providing computing resources for the HPC System, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. We appreciate helpful conversations with Prof. Emanuela Del Gado, Prof. Silke Henkes, Prof. Heinrich Jaeger, and Dr. Mike van der Naald, to the last of whom we are grateful for assistance in implementing the pebble game algorithm.

### v5

@@ Acknowledgments
This work was supported by NSF CBET-2228681, CBET-2228680, and DMR-2026834. AS acknowledges Aryabhata HPC facility of IIT (ISM) Dhanbad and National Supercomputing Mission (NSM) for computing resources on the HPC System, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. We appreciate helpful conversations with Emanuela Del Gado, Silke Henkes, Heinrich Jaeger, and Mike van der Naald, to the last of whom we are grateful for assistance in implementing the pebble game algorithm.